OpenStudy (anonymous):

how to get the formula for number of classes and class interval in stat>?? =D the given is: 101,78,75,76,65,76,57,60,82,68,76,75,83,79,42,72,87,,80,76,80

OpenStudy (mathmate):

If there are no outliers, we usually have the number of classes equal to k=sqrt(n), where n = number of data items. The class interval with then be (range+1)/k, rounded to the next convenient number.

OpenStudy (anonymous):

ahahah.. can't freaking undertand. =D i am still in 2nd year dude. =D we havn't got to that k=sqrt(n),

OpenStudy (anonymous):

i already got the range which is 7, but i can't seem to get the starting point. =D

OpenStudy (anonymous):

how to get the starting point??

OpenStudy (mathmate):

The sqrt(n) thing means that if you have 100 data items, you would make sqrt(100)=10 classes, and so on. To find the class starting point, you need to find the smallest number S and the largest number L. The difference L-S gives you the range. We usually start a class at S-1 or S. The class size is (L-S)/k, rounded to a convenient number. In your case, L=101, S=42. so the range is 101-42=59. n=20, k=sqrt(20)=4.5 =>5 Class size = range/k ...

OpenStudy (anonymous):

tnx, =D

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

oh.. hahaha sqrt is square root. hahah =D

OpenStudy (mathmate):

Sorry, I didn't explain that!

OpenStudy (anonymous):

i thought class size is supposed to be : range/number of classes then round up.

OpenStudy (mathmate):

Yes, you're right. I put k=sqrt(n), but it shoud be k=sqrt(n) rounded up or something.

OpenStudy (anonymous):

tnx,, again. haha bye

OpenStudy (mathmate):

you're welcome! :)

OpenStudy (anonymous):

where are you from ???

OpenStudy (mathmate):

Canada